OpenStudy (anonymous):

Solve the system of equations by the addition method. x+y=-12 x-y=-12

OpenStudy (anonymous):

If x + y = 12, and x - y = -12, y needs to be zero, correct? That would mean x is -12... Am I missing something here? Is there more information to this question?

OpenStudy (anonymous):

they both equal to -12.

OpenStudy (anonymous):

That's the only information/instructions.

OpenStudy (anonymous):

With that being said, x = -12, and y = 0, but that's way too easy. There must be something I'm missing.

OpenStudy (anonymous):

its a multiple choice question.. o:

OpenStudy (anonymous):

What are the options?

OpenStudy (anonymous):

(0, 12) (12, 0) (0, -12) (-12, 0)

OpenStudy (anonymous):

Could you like show the step by step process of the work? O:

OpenStudy (anonymous):

(x + y) & (x - y) = -12 y, as aforementioned, has to be zero, for the outcome remains the same if both subtracted and added. (x + 0) & (x - 0) = -12 For addition: x + y = -12 -12 - y = x -12 - 0 = x -12 - 0 = -12 -12 = x For subtraction, flip the signs.

OpenStudy (anonymous):

oh. I understand. Thank you ^^;;

OpenStudy (anonymous):

No problem. Any other questions you need help with?

OpenStudy (anonymous):

Most of the questions on this quiz I am taking on blackboard is all related with systematic equations. . .

OpenStudy (anonymous):

I can help.

OpenStudy (anonymous):

8x+4y=-8 -2x-4y=-28

OpenStudy (anonymous):

Multiple choice answers: (4, -10) (8, -10) (-6, 10) infinite number of solutions

OpenStudy (anonymous):

I just want to learn how to approach this type of problem.

OpenStudy (anonymous):

4y is a shared sequence, so finding that in one equation can potentially solve both... I'm working on it.

OpenStudy (anonymous):

ok o :